OpenStudy (kaylala):

proving identities: (tan^2+1)/(tan+cot) = tan

OpenStudy (kaylala):

@superdavesuper

OpenStudy (superdavesuper):

use same eqn like last time sin^2 + cos^2 = 1 to simplify n solve but in this case, replace tan=sin/cos and cot=cos/sin first then multiply top n bottom by sin*cos.... u will need to go thro a few steps...list them out here n we will see

OpenStudy (kaylala):

\[((\sin \div \cos ) + 1 ) \div ((\sin \div \cos ) + (\cos \div \sin ))\]

OpenStudy (kaylala):

?

OpenStudy (superdavesuper):

yup yup

OpenStudy (kaylala):

what's next?

OpenStudy (superdavesuper):

\[[((\sin ^{2}x/\cos ^{2}x) + 1)*(sinxcosx)]/[(sinx/cosx + cosx/sinx)*(sinxcosx)]\]

OpenStudy (kaylala):

will it be : (sin^3 + cos^2sin) / (sin^2+cos^2)

OpenStudy (superdavesuper):

scroll up to see the 1st eqn i wrote n the bottom can be simplified to....

OpenStudy (kaylala):

i don't get the "n" part what's that supposed to mean?

OpenStudy (superdavesuper):

sorry sorry....i mean to say scroll up....AND use that eqn to simplify the bottom

OpenStudy (kaylala):

only sin gets left

OpenStudy (kaylala):

i cancelled out the rest only sin gets left @superdavesuper

OpenStudy (superdavesuper):

well that is because ur top part was not right....please try to expand that again :)

OpenStudy (kaylala):

i see now. thanks!

OpenStudy (superdavesuper):

welcome :)